OpenStudy (anonymous):

a group purchased 10 hotdogs and 5 drinks at a cost of 320. a second group bought 7 hotdogs and 4 drinks at 228.50. how much is the cost of one hotdog? a drink?

OpenStudy (nincompoop):

set things up as a systems of equation then solve for the hotdog

OpenStudy (nincompoop):

that is some funny looking letter h

OpenStudy (nincompoop):

@punyeta andyan ka?

OpenStudy (anonymous):

wla n, tulog n...XD

OpenStudy (anonymous):

it is a giraffe looking back

OpenStudy (nincompoop):

let me give you an example 5x + 3y = 18 (1) 2x - 6y = 30 (2) pick what you want to solve first, let us say y. multiply the first equation by 2 so it becomes 10x + 6y = 36 which means our y can be canceled 2x - 6y = 30 ---------------- 12x = 66 divide both sides by 12 x = 11/5 use the value of x and plug it in to one of the original equation to solve for y 2(11/5) - 6y = 30 22/5 - 6y = 30 -6y = 30- 22/5 simplify

OpenStudy (nincompoop):

anymore questions?